OpenStudy (anonymous):

Can someone explain this for me step by step? Abby rows 10 km upstream and 10 km back in a total time of 3 hr. The speed of the river is 5 km/h. Find Abby’s speed in still water.

OpenStudy (anonymous):

am i using r=d/t

hero (hero):

use d = rt and let r = b ± w

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

what does let r = b ± w mean

hero (hero):

t1 = Upstream time t2 = Downsteam time 10 = 5t 2 = t1 t1 + t2 = 3 2 + t2 = 3 t2 = 3 - 2 t2 = 1 I'll show you what that means in a sec

OpenStudy (anonymous):

ok thank you

hero (hero):

Upstream Formula d = rt1 d = (b - w)t1 Downsteam Formula: d = rt2 d = (b + w)t2

hero (hero):

b - w = 5

hero (hero):

Upstream 10 = 2(b - w) Downsteam 10 = (b + w) 10 = 2b - 2w 20 = 2b +2w 30 = 4b 30/4 = b 7.5 = b

hero (hero):

10 = b + w 10 = 7.5 + w 10 - 7.5 = w 2.5 = w

hero (hero):

Therefore The speed of the boat was 7.5 km/h

hero (hero):

Any questions?

OpenStudy (anonymous):

do you have to do both equations

hero (hero):

You don't have to solve for w in this case, however, it is good to practice doing so.

OpenStudy (anonymous):

ok that makes it easier to follow

OpenStudy (anonymous):

so if I put both equations then I would be correct as well

OpenStudy (anonymous):

what does the b and w stand for if you dont mind

OpenStudy (anonymous):

hello

OpenStudy (anonymous):

so confused

hero (hero):

b = speed of the boat w = speed of the water current

hero (hero):

@andjie

OpenStudy (anonymous):

Thank you

hero (hero):

All you have to do is just keep doing similar problems like these. You'll get the hang of it.

OpenStudy (anonymous):

thank you so much, sorry I left